If an amplifier working with ±5V DC power supplies, requires a 

minimum output voltage of 3V across a load resistance of 100 ohms, 

calculate the minimum current to be drawn from the power supply, if 

the efficiency of the amplifier is 35%.


1
Expert's answer
2021-02-25T15:04:30-0500

"\eta=\frac{U_mI_m}{U_mI_m+\frac{U^2}{R}},\implies" "I_m=\frac{\eta U^2}{(1-\eta)U_mR},"

"I_m=\frac{0.35\cdot 5^2}{0.65\cdot 3\cdot 100}=44.9~mA."